OpenStudy (anonymous):

Solve the following separable ODE y^1=sec^2y

OpenStudy (freckles):

\[y'=\sec^2(y) \\ \frac{dy}{dx}=\sec^2(y) \\ \frac{dy}{\sec^2(y)}=dx \\ \cos^2(y) dy=dx\] integrate both sides

OpenStudy (anonymous):

Solve the following separable ODE xy^1 + y = 0, y(4) = 6

OpenStudy (freckles):

separate the variables again subtract y on both sides then divide both sides by x and y and multiply dx on both sides